i m having problem in learning the basic formulas and properties of conics. can some1 sujjest me ways to learn them??

Grade:12

3 Answers

Sathya
12 years ago

Hi Manu Lahariya,

 

The Conic Sections. For any of the below with a center (j, k) instead of (0, 0), replace each x term with (x-j) and each y term with (y-k).

  Circle Ellipse Parabola Hyperbola
Equation (horiz. vertex): x2 + y2 = r2 x2 / a2 + y2 / b2 = 1 4px = y2 x2 / a2 - y2 / b2 = 1
Equations of Asymptotes:       y = ± (b/a)x
Equation (vert. vertex): x2 + y2 = r2 y2 / a2 + x2 / b2 = 1 4py = x2 y2 / a2 - x2 / b2 = 1
Equations of Asymptotes:       x = ± (b/a)y
Variables: r = circle radius a = major radius (= 1/2 length major axis)
b = minor radius (= 1/2 length minor axis)
c = distance center to focus
p = distance from vertex to focus (or directrix) a = 1/2 length major axis
b = 1/2 length minor axis
c = distance center to focus
Eccentricity: 0   c/a c/a
Relation to Focus: p = 0 a2 - b2 = c2 p = p a2 + b2 = c2
Definition: is the locus of all points which meet the condition... distance to the origin is constant sum of distances to each focus is constant distance to focus = distance to directrix difference between distances to each foci is constant

 

Pls approve my answer if it was helpful Wink

Vinayak Gupta
12 years ago

I WOULD SUGGEST YOU TO FIRST UNDERSTAND THE BASIC DIFFERENCE BETWEEN THE CONICS AND THEN THEIR GRAPHS AS THE MAIN DIFFERENCE BETWEEN THEM IS VISIBLE CLEARLY ON THE GRAPHS.

 

THEN,

I WOULD SUGGEST YOU LOOK AT THE STUDY MATERIAL PROVIDED BY YOUR TUTIONS/COACHING AS THEY GIVE A BRIEF IDEA IN A RATHER SHORT AND PRECISE WAY.

THIS IS PROVIDED IN A VERY GOOD WAY IN THE ASKIITIANS WEBSITE.

THERE IS NO NEED TO MUG UP EACH FORMULAE BECAUSE QUESTIONS IN IIT JEE AND AIEEE REQUIRE A HIGH LEVEL UNDERSTANDING OF EACH CONCEPT AND THEREFORE IF YOU ARE NOT ABLE TO UNDERSTAND EACH CONCEPT CLEARLY YOU WOULD NOT BE ABLE TO SOLVE THE QUESTIONS AND MAKE SOME MISTAKES.

GOOD LUCK

annie mughal
12 years ago

solve as many questions as you can. first some while looking at those formulas and the try doing some more without looking at them! hide them and then practise! memorize them!